# Enhanced Smart Contract Reputability Analysis using Multimodal Data Fusion on Ethereum

### v1

@@ VI Concluding Remarks
This study demonstrated the effectiveness of leveraging static code analysis for smart contract reputability prediction, showing that opcode embeddings provide valuable insights into contract vulnerabilities and reputability trends. By employing boosting algorithms and anomaly detection techniques, the framework accurately identified illicit contracts, reinforcing the role of code-based features in distinguishing reputable behaviour. However, static features alone struggle to capture evolving behavioural shifts, highlighting the need for a more comprehensive approach that considers dynamic interactions and real-time patterns.
To address this, a multimodal data fusion framework was developed, integrating transactional data with static code features to capture both structural and behavioural anomalies. The inclusion of transactional data significantly enhanced the model’s ability to distinguish true anomalies from benign deviations, demonstrating the value of combining static and dynamic information for a more holistic reputability assessment. This capability is particularly relevant for DeFi platforms, where smart contracts facilitate critical processes such as lending, staking, and token exchanges, often handling substantial financial transactions autonomously. In such high-stakes environments, the early detection of anomalous contract behaviour is crucial for preserving user trust and minimizing potential financial losses. The ability to detect illicit activity in near real-time not only strengthens the resilience of DeFi ecosystems but also supports regulatory compliance efforts by providing actionable insights into contract performance and reputability trends.

### v2

@@ VI Concluding Remarks
This study demonstrated the effectiveness of leveraging AI-based code analysis for smart contract reputability prediction, showing that opcode embeddings provide valuable insights into contract vulnerabilities and reputability trends. By employing boosting algorithms and anomaly detection techniques, the framework accurately identified illicit contracts, reinforcing the role of code-based features in distinguishing reputable behaviour. However, code features alone struggle to capture evolving behavioural shifts, highlighting the need for a more comprehensive approach that considers dynamic interactions and real-time patterns.
To address this, a multimodal data fusion framework was developed, integrating transactional data with code features to capture both structural and behavioural anomalies. The inclusion of transactional data significantly enhanced the model’s ability to distinguish true anomalies from benign deviations, demonstrating the value of combining code and dynamic information for a more holistic reputability assessment. This capability is particularly relevant for DeFi platforms, where smart contracts facilitate critical processes such as lending, staking, and token exchanges, often handling substantial financial transactions autonomously. In such high-stakes environments, the early detection of anomalous contract behaviour is crucial for preserving user trust and minimizing potential financial losses. The ability to detect illicit activity in near real-time not only strengthens the resilience of DeFi ecosystems but also supports regulatory compliance efforts by providing actionable insights into contract performance and reputability trends.